Three moles of a gas expand isothermally. If the initial temperature of the gas is 300K and its final volume is twice the initial volume, calculate the work done by the gas.
1
Expert's answer
2020-12-06T17:28:08-0500

"W=\intop pdV=\nu RT\int \frac{dV}{V}=\nu RT \ln{\frac{V_2}{V_1}}="


"=3\cdot 8.31\cdot 300 \cdot\ln 2=5184(J)"